OpenStudy (anonymous):

In how many ways can a committee of five men and four women be formed from a group of nine men and nine women?

OpenStudy (anonymous):

\[\dbinom{9}{5}\times \dbinom{9}{4}\] do you know how to compute these numbers?

OpenStudy (anonymous):

im using permutations and combinations though

OpenStudy (anonymous):

is that still the right answer?

OpenStudy (anonymous):

it is 9 choose 5 times 9 choose 4

OpenStudy (anonymous):

maybe you have seen \[\dbinom{9}{5}\] written as \[_9C_5\] it is the same thing

OpenStudy (anonymous):

soo on the calculator would it be 9 C 5

OpenStudy (anonymous):

oh! okay sweet

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and in fact 9 choose 5 is the same number as 9 choose 4

OpenStudy (anonymous):

okay and then when you get the two seperate answers do you divide?

OpenStudy (anonymous):

oh no you multiply!

OpenStudy (anonymous):

ohh! okay

OpenStudy (anonymous):

you get \[\dbinom{9}{5}=\dbinom{9}{4}=126\] so your answer is \[126\times 126=126^2\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw